OpenStudy (anonymous):

express in its simplest form: 5x-6/x²+x-6 - 3/x+3

OpenStudy (anonymous):

write it in proper way

OpenStudy (anonymous):

\[\left(\begin{matrix}5x-6 \\ x²+x-6\end{matrix}\right)-\left(\begin{matrix}3 \\ x+3\end{matrix}\right)\]

OpenStudy (anonymous):

dont know how to get the division line soz, ignore the brackets aswell =/

OpenStudy (anonymous):

\[\frac{5x-6}{x^2+x-6}-\frac{3}{x+3}\]

OpenStudy (anonymous):

ya thats it

OpenStudy (anonymous):

factor first fraction's denominator

OpenStudy (anonymous):

(x-2)(x+3)

OpenStudy (anonymous):

....what do i do now?

OpenStudy (anonymous):